OpenStudy (anonymous):

if A and B are independent events and P(A) and P(B) are both greater than 1/2, than P(A and B) is ______ greater than 1 A. Sometimes B always C never

OpenStudy (anonymous):

permutation

OpenStudy (kirbykirby):

If A and B are independent, then \(P(A \text{ and } B )=P(A \cap B) = P(A)P(B)\) So a fraction times a fraction (less than 1) is always less than 1. Besides, a probability can never be greater than 1 anyway. Even if P(A) = 0.99 and P(B) = 0.99, then P( A and B) = P(A)P(B) = 0.9801.

OpenStudy (anonymous):

So is it never